OpenStudy (anonymous):

solve using the multiplication principle -1/2x = -5/6

OpenStudy (anonymous):

can someone check my answer plse solve using the multiplication principle -1/2x = -5/6 x=3/5

OpenStudy (anonymous):

if x is in denominator your answer is correct.elseif it is in numerator x=5/3

OpenStudy (anonymous):

u have to multiply

OpenStudy (anonymous):

george7 x is placed in the same spot as the underline

OpenStudy (anonymous):

rich kidz how can you tell if x is in denominator or numerator I see the x is placed in the same spot as the underline so what is it

OpenStudy (anonymous):

numerator

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

rich kidz so the answer would be x=5/3

OpenStudy (anonymous):

am not real smart

OpenStudy (anonymous):

yeah that IT